# WinkTPG: An Execution Framework for Multi-Agent Path Finding Using Temporal Reasoning

### v1

@@ 3.4 Satisfying Conflicting Edges
5 Windowed kTPG (WinkTPG)
While adding safety margins ensures the speed profiles are collision-free with a certain probability, as shown in Eq. 2, the uncertainty in the reach times of an agent increases along its path. So, according to Eq. 4, the required safety margin for Type-2 edges between later vertices may become large, compromising solution quality. To address this, we propose using online execution information to reduce uncertainty.
Definition 3 (Real-Time MAEP under Uncertainty)
We extend the MAEP with uncertainty in Definition 2 by including information updates during execution, as proposed in (Hönig et al. 2019). We assume that agents report their actual reach times and current kinodynamic states upon arriving at each location.
To maintain robustness while improving solution quality, we propose WinkTPG. WinkTPG periodically replaces the estimated reach times with the actual reach times reported by agents. Using these actual times, it recalculates the estimated reach times for subsequent vertices, reducing their uncertainty. WinkTPG employs a windowed replanning method (Berndt et al. 2023) to incorporate this updated information. Concretely, we trigger replanning every $t_{e}$ seconds, during which we invoke kTPGu to generate speed profiles within a defined planning window. A planning window, shown as the green region in Fig. 5, is a subset of the original TPG and spans at least $t_{p}$ vertices for each agent.
When replanning is triggered, we follow the method from (Berndt et al. 2023) to determine which vertices to include in the planning window. For each agent, we identify its most recently reported vertex and the corresponding reach time. The next $N_{E}$ vertices are then marked as enqueued, meaning the agent can continue to follow its previous speed profile through these vertices without replanning. The planning window begins immediately after the last enqueued vertex. As shown in Fig. 5, we first include the next $t_{p}$ vertices for each agent in the window. To ensure all precedence constraints are satisfied, we further check Type-2 edges within the window. If any such edge originates from outside the current window (e.g., the red edge in Fig. 5), we recursively expand the window to include the source vertex and any earlier vertices needed to preserve execution order.
6 Theoretical Analysis
Theorem 1 (Completeness of kTPG and kTPGu)
Given a collision-free MAPF plan, kTPG and kTPGu return a feasible solution in finite time, though it may be suboptimal.
Proof 1
As shown in Section 3.4, our method of determining the split time ensures that we can always find a valid speed profile in each iteration. Since $\mathcal{G}$ is guaranteed to be a directed acyclic graph (Hönig et al. 2016), if there are locked vertices in $\mathcal{G}$, there is at least one conflicting edge originating from an unlocked vertex to a locked vertex. So when we replan for the agent associated with this unlocked vertex, kTPG guarantees progress by satisfying at least this conflicting edge. Therefore, kTPG is guaranteed to unlock all vertices after a finite number of iterations. That is, kTPG is complete and guarantees to find feasible solutions. For kTPGu, incorporating a safety margin does not affect the existence of a valid speed profile in each iteration, making the completeness of kTPG unchanged.
Theorem 2 (Completeness of WinkTPG)
Given a collision-free MAPF plan, WinkTPG ensures that all agents reach their goal locations without collisions in finite time.
Proof 2
In each planning window, feasibility is ensured by allowing each agent to retain and follow its previously computed speed profile up to the last vertex from the previous window. Based on the reasoning in Section 3.4, we ensure that, for each agent, a valid speed profile exists for each agent in the current window, enabling it to wait at the final vertex of the previous window before proceeding. As shown in Theorem 1, at least one agent makes progress in each set of speed profiles. Thus, WinkTPG completes in finite time with a bounded number of kTPGu calls.
Theorem 3 (Time Complexity of kTPG and kTPGu)
Proof 3
As proved in Theorem 1, each iteration satisfies at least one new Type-2 edge and unlocks at least one previously locked vertex. This implies that the total number of iterations is upper bounded by $\min(|\mathcal{V}|,|\mathcal{E}_{2}|)$. Since WinkTPG calls kTPGu within each of the $W$ planning windows, the total number of iterations is at most $W\cdot\min(|\mathcal{V}|,|\mathcal{E}_{2}|)$.
7 Empirical Evaluation
All methods included in the study were implemented in C++ and tested on an Ubuntu 20.04 system with an AMD 3990x processor and 188 GB of memory. The code was run using a single core for all computations. The source code for our method will be publicly available upon acceptance.
@@ 7 Empirical Evaluation
We evaluate all methods on six four-neighbor grid maps from the MovingAI MAPF benchmark (Stern et al. 2019): empty (size: 32$\times$32), random (size: 64$\times$64), room (size: 64$\times$64), den520d (size: 256$\times$257), and Boston (size: 256$\times$256), warehouse-large (size: 340×164340$\times$ 164340 × 164). The distance between two neighboring grids is $1\;m$. The MAPF plans used for execution are generated by PBS (Ma et al. 2019). We use suboptimality ratio to evaluate the solution quality, which is defined as $\frac{T_{sum}-T_{ideal}}{T_{ideal}}$, where $T_{sum}$ is the sum of reach times, and $T_{ideal}$ is the sum of individual agent reach times without considering collisions or delays.
7.1 Kinodynamic Planning
@@
In this experiment, we evaluate the solution quality of the speed profiles generated by kTPG and compare it against two baselines: ADG and KDN. ADG (Hönig et al. 2019) generates speed profiles by planning only for vertices without unsatisfied incoming Type-2 edges, using the same speed profile planner as kTPG. Furthermore, ADG is assumed to have access to real-time agent locations and replans its speed profile every $0.01s$. KDN (Zhang et al. 2021) formulates the temporal dependency of TPG as a MILP problem and finds the optimal solution for Definition 1.
@@
[…]
@@
As shown in Fig. 6, both kTPG and KDN show significant improvements in solution quality across all maps compared to ADG. This improvement is more evident on larger maps with more agents (e.g., up to 51.7% in the warehouse map). This is because ADG only plans speed profiles for vertices without incoming Type-2 edges. In situations with more Type-2 edges, this strategy is overly conservative and leads to frequent unnecessary deceleration. While KDN achieves slightly better solution quality, it is much less scalable than kTPG. Both kTPG and ADG maintain a 100% success rate on problems involving up to 1,000 agents. In contrast, KDN, while ensuring optimality, cannot solve problems with more than 100 agents. This is due to KDN relying on a MILP formulation to encode the precedence constraints imposed by Type-2 edges, whose complexity grows exponentially with the number of Type-2 edges. Notably, while the runtime for each planning round of ADG is low, its frequent replanning significantly increases the total runtime, making it similar to that of kTPG.
Figure 7: Comparison of solution cost between ADG (dashed lines) and WinkTPG (solid lines) across different uncertainty levels (indicated by color). The ADG lines largely overlap across uncertainty levels in both figures.
Figure 8: Comparison of solution cost between ADG and WinkTPG, with different execution and planning window.
7.2 Robust Execution Under Uncertainty
In this section, we evaluate the performance of WinkTPG in environments with temporal uncertainty. As KDN cannot handle uncertainty, only ADG is used as a baseline.
@@ 7.2 Robust Execution Under Uncertainty
[…]
Result Analysis
[…]
@@
Next, we evaluate the relationship between solution quality and the length of the execution and planning windows. All experiments are conducted in environments with $\varepsilon=0.03$. As shown in Fig. 8, configurations where the execution window is smaller than the planning window consistently show better solution quality. Also, smaller execution windows always achieve better performance, due to more frequent replanning with updated location information.
@@
Table 1: Planning Runtime (in seconds) on the warehouse map with $\varepsilon=0.03$. MAPF refers to the time taken to generate the initial MAPF plan, while ADG and WinkTPG represent the runtime per planning round.
Agents # MAPF ADG WinkTPG
100 0.089±0.030.089$\pm$ 0.030.089 ± 0.03
2.36±0.292.36$\pm$ 0.292.36 ± 0.29
×$\times 10^{-4}$
$t_{p}$=10 s 0.002±0.000.002$\pm$ 0.000.002 ± 0.00
$t_{p}$=20 s 0.003±0.000.003$\pm$ 0.000.003 ± 0.00
$t_{p}$=40 s 0.005±0.000.005$\pm$ 0.000.005 ± 0.00
$t_{p}$$\infty$ 0.266±0.050.266$\pm$ 0.050.266 ± 0.05
1,000 44.51±3.9444.51$\pm$ 3.9444.51 ± 3.94
2.02±0.092.02$\pm$ 0.092.02 ± 0.09
×$\times 10^{-3}$
$t_{p}$=10 s 0.344±0.010.344$\pm$ 0.010.344 ± 0.01
$t_{p}$=20 s 0.389±0.010.389$\pm$ 0.010.389 ± 0.01
$t_{p}$=40 s 0.519±0.020.519$\pm$ 0.020.519 ± 0.02
$t_{p}$$\infty$ 28.21±1.6728.21$\pm$ 1.6728.21 ± 1.67
@@
Finally, we report in Table 1 the runtime of ADG, WinkTPG, and the MAPF planner. In both cases, ADG and WinkTPG demonstrate significantly faster runtimes compared to the MAPF planner used to generate the paths. Notably, while ADG appears faster, it requires significantly more frequent replanning. Additionally, with an appropriately chosen planning window size, WinkTPG generates plans in under a second, showing it is capable of near real-time performance.
8 Conclusion

### v2

@@ IV-C Integration with kTPG
V Windowed kTPG (WinkTPG)
While adding safety margins ensures the speed profiles are collision-free with a certain probability, as shown in Eqs. 6, 8 and 10, the uncertainty in the reach times of an agent increases along its path. So, according to Eq. 11, the required safety margin for Type-2 edges between later vertices may become large, compromising solution quality. To address this, we propose using online execution information to reduce estimated execution timing uncertainty.
V-A Real-Time MAEP under Uncertainty
We extend the MAEP with uncertainty in Section IV-A by including information updates during execution, as proposed in [9]. We assume that agents report their actual reach times and current kinodynamic states upon arriving at each location.
V-B WinkTPG
To maintain robustness while improving solution quality, we propose WinkTPG. WinkTPG replaces the estimated reach times with the actual reach times reported by agents. Using these actual times, it recalculates the estimated reach times for subsequent vertices, reducing their uncertainty. WinkTPG employs a windowed replanning method [5] to incorporate this updated information. Concretely, we trigger replanning every $t_{e}$ seconds, during which we invoke kTPGu to generate speed profiles within a defined planning window. A planning window, shown as the green region in Fig. 6, is a subset of the original TPG and spans at least $t_{p}$ vertices for each agent.
When replanning is triggered, we follow the method from [5] to determine which vertices to include in the planning window. For each agent, we identify its most recently reported vertex and the corresponding reach time. The next $N_{E}$ vertices, with $N_{E}$ as a predefined hyper-parameter, are then marked as enqueued, meaning the agent can continue to follow its previous speed profile through these vertices without replanning. The planning window begins immediately after the last enqueued vertex. As shown in Fig. 6, we first include the next $t_{p}$ vertices for each agent in the window. To ensure all precedence constraints are satisfied, we further check Type-2 edges within the window. If any such edge originates from outside the current window (e.g., the orange edge in Fig. 6), we recursively expand the window to include the source vertex and any earlier vertices needed to preserve execution order.
VI Theoretical Analysis
Theorem 1 (Completeness of kTPG and kTPGu).
Given a collision-free and cycle-conflict-free MAPF plan, kTPG and kTPGu terminate in finite time and produce a kinodynamically feasible solution that preserves all precedence constraints.
Proof sketch.
We first establish the existence of a feasible solution by leveraging proof from [10, 9]. As shown in these works, if the original MAPF plan is collision-free and cycle-conflict-free, then a kinodynamically feasible set of speed profile that preserves all precedence constraints exists. We now prove termination of kTPG. As shown in Section III-E, our method of determining the split time ensures that we can always find a valid speed profile in each iteration. Since $\mathcal{G}$ is guaranteed to be a directed acyclic graph, if there are locked vertices in $\mathcal{G}$, there is at least one conflicting edge originating from an unlocked vertex to a locked vertex. So when we replan for the agent associated with this unlocked vertex, kTPG guarantees progress by satisfying at least this conflicting edge. Therefore, kTPG is guaranteed to unlock all vertices after a finite number of iterations. That is, kTPG is complete and guarantees to find feasible solutions. For kTPGu, incorporating a safety margin does not affect the existence of a valid speed profile in each iteration, making the completeness of kTPG unchanged. ∎
Theorem 2 (Completeness of WinkTPG).
Given a collision-free and cycle-conflict-free MAPF plan, WinkTPG ensures that, under the adopted execution timing uncertainty model, all agents reach their goal locations in finite time, while precedence satisfaction and collision-free execution are preserved according to the guarantees provided by that model.
Proof sketch.
In each planning window, feasibility is ensured by allowing each agent to retain and follow its previously computed speed profile up to the last vertex from the previous window. Based on the reasoning in Section III-E, we ensure that, for each agent, a valid speed profile exists in the current window, enabling it to wait at the final vertex of the previous window before proceeding. As shown in Theorem 1, at least one agent makes progress in each set of speed profiles. Thus, WinkTPG completes in finite time with a bounded number of kTPGu calls. ∎
Theorem 3 (Time Complexity of kTPG and kTPGu).
Proof.
As proved in Theorem 1, each iteration satisfies at least one new Type-2 edge and unlocks at least one previously locked vertex. This implies that the total number of iterations is upper bounded by $\min(|\mathcal{V}|,|\mathcal{E}_{2}|)$. ∎
Since WinkTPG calls kTPGu within each of the $W$ planning windows, the total number of iterations is at most $W\cdot\min(|\mathcal{V}|,|\mathcal{E}_{2}|)$.
VII Empirical Evaluation
@@
In this section, we empirically evaluate the proposed methods. First, we compare kTPG with prior MAPF execution methods in terms of solution quality and scalability under kinodynamic constraints. Second, we study WinkTPG under different execution timing uncertainty models. Third, we validate WinkTPG in more realistic settings using both physics-based simulation and physical-robot experiments.
@@
We evaluate all methods on six maps with 25 instances from the MovingAI MAPF benchmark [26]: empty (empty-32-32), random (random-64-64-10), room (room-64-64-16), den520d (den520d), and Boston (Boston_0_256), warehouse (warehouse-20-40-10-2-2). The distance between two neighboring cells is 1m1\;m. The MAPF plans used for execution are generated by PBS [18]. We use suboptimality ratio to evaluate the solution quality, which is defined as $\frac{T_{sum}-T_{ideal}}{T_{ideal}}$, where $T_{sum}$ is the sum of reach times, and $T_{ideal}$ is the sum of individual agent reach times without considering collisions or delays.
Unless otherwise specified, all methods included in the study were implemented in C++ and tested on an Ubuntu 20.04 system with an AMD 3990x processor and 188 GB of memory. The code was run using a single core for all computations. The source code for our method is publicly available at https://github.com/JingtianYan/IEEE-T-ASE-WinkTPG.
VII-A Kinodynamic Planning
@@
In this experiment, we evaluate the solution quality of the speed profiles generated by kTPG and compare it against two baselines: ADG and KDN. ADG [9] generates speed profiles by planning only for vertices without unsatisfied incoming Type-2 edges, using the same speed profile planner as kTPG. Furthermore, ADG is assumed to have access to real-time agent locations and replans its speed profile every 0.01s0.01\,s. KDN [32] formulates the temporal dependency of TPG as a MILP problem and finds the optimal solution for MAEP (see Section III-A).
@@
Since KDN is designed for an omnidirectional robot model without uncertainty, we adopt the same setting in this experiment. An omnidirectional robot can move in any orientation, with its speed profile constrained by the dynamic limits of its physical model. In this experiment, each agent is subject to a speed limit of $0,2]\,\text{m/s}$ and an acceleration limit of $1,1]\,\text{m/s}^{2}$. As required by the speed profile planner, speeds at the vertices are discretized into three levels: $\{0,\sqrt{2},2\}\,\text{m/s}$. The runtime limit for all methods is 300 seconds.
@@
As shown in Fig. 7, both kTPG and KDN show significant improvements in solution quality across all maps compared to ADG. This improvement is more evident on larger maps with more agents (e.g., up to 51.7% on the warehouse map). This is because ADG only plans speed profiles for vertices without incoming Type-2 edges. In situations with more Type-2 edges, this strategy is overly conservative and leads to frequent unnecessary deceleration. While KDN achieves slightly better solution quality, it is much less scalable than kTPG. Both kTPG and ADG maintain a 100% success rate on problems involving up to 1,000 agents. In contrast, KDN, while ensuring optimality, cannot solve problems with more than 200 agents. This is due to KDN relying on a MILP formulation to encode the precedence constraints imposed by Type-2 edges, whose complexity grows exponentially with the number of Type-2 edges. Notably, while the runtime for each planning round of ADG is low, its frequent replanning significantly increases the total runtime, making it similar to that of kTPG.
TABLE I: Additional reliability and efficiency metrics on the warehouse map with 100 agents under three execution timing uncertainty models ($t_{e}=10$, $t_{p}=20$). Results are averaged over all scenarios.
Noise level Violation rate (%) Avg. velocity (m/s) Total idle time (s) Total stop-and-go count
ADG WinkTPG ADG WinkTPG ADG WinkTPG ADG WinkTPG
$\sigma=0.001$ 0.00 0.74 0.95 1.02 3460.00 3104.22 3920.88 3317.72
$\sigma=0.03$ 0.00 0.70 0.95 1.01 3481.46 3236.12 3920.52 3320.28
$\sigma=0.05$ 0.00 0.69 0.95 1.01 3497.71 3330.00 3919.40 3322.28
$\mathcal{K}=1.0$ 0.00 0.00 0.95 1.02 3453.82 3164.25 3920.48 3324.00
$\mathcal{K}=3.0$ 0.00 0.00 0.94 1.01 3504.14 3227.75 3922.76 3327.24
$\mathcal{K}=5.0$ 0.00 0.00 0.93 0.99 3544.29 3305.45 3923.52 3328.58
$\epsilon=0.01$ 0.00 0.00 0.95 1.02 3461.88 3147.40 3920.08 3320.24
$\epsilon=0.02$ 0.00 0.00 0.95 1.02 3468.71 3196.95 3922.12 3323.28
$\epsilon=0.1$ 0.00 0.00 0.95 0.99 3506.42 3616.23 3920.04 3345.20
VII-B Robust Execution Under Uncertainty
In this section, we evaluate the performance of WinkTPG in environments with execution timing uncertainty. As KDN cannot handle uncertainty, only ADG is used as a baseline. Our goal is to examine whether the proposed safety margin mechanism can provide the same guarantees predicted by the theory. We also study its execution efficiency under different factors, including window size and uncertainty level.
Figure 8: Comparison of solution quality between ADG (solid lines) and WinkTPG (dashed lines) under the Gaussian noise model across different uncertainty levels (indicated by color). The ADG lines largely overlap across uncertainty levels in both figures. Shaded regions indicate the variance across all scenarios.
Figure 9: Comparison of solution quality between ADG and WinkTPG under the Gaussian noise model, with different execution and planning windows.
@@ VII-B Robust Execution Under Uncertainty
We use a differential-drive robot model that is more commonly used in real-world applications. Compared to omnidirectional robots, it moves only along its current orientation, while being constrained by the same speed and acceleration limits as in the previous section. It can change its orientation only when its velocity is zero. Specifically, it takes $0.5\,\text{s}$ to turn left or right and $0.9\,\text{s}$ to turn backward. To evaluate both reliability and execution efficiency, we report the following additional metrics in Table I: (1) Precedence violation rate: the ratio of violated precedence constraints at shared locations to the total number of such constraints. (2) Average velocity: the mean linear velocity over the entire execution horizon. (3) Total idle time: the cumulative duration during which the linear and angular velocity of an agent is zero. (4) Stop-and-go count: the number of times an agent accelerates from zero velocity to positive velocity.
VII-B1 Unbounded Uncertainty Model
We first evaluate WinkTPG under a Gaussian noise model. We parameterize the noise magnitude using a scalar $\sigma$ shared across all agents By Eq. 1, this implies that the traversal time between adjacent vertices (distance $1\,\text{m}$) has variance $\sigma^{2}$. For intuition, the nominal traversal time between adjacent vertices at maximum speed is $\frac{1\,\text{m}}{2\,\text{m/s}}=0.5\,\text{s}$. When $\sigma=0.05$, applying the $3\sigma$ rule [6] implies up to $\pm 30\%\times 0.5\,\text{s}$ deviation. Since this already represents substantial timing variability for short-range motion, we restrict $\sigma\leq 0.05$. We set the safety probability threshold to $P_{d}=0.95$ unless otherwise specified.
Effect of Uncertainty Magnitude. We first evaluate WinkTPG under varying uncertainty levels using a fixed window size. Specifically, the execution window $t_{e}$ is set to $10\,\text{s}$, and the planning window $t_{p}$ is $20$. As shown in Fig. 8, while WinkTPG generates more suboptimal solutions with larger $\sigma$, it consistently outperforms ADG in environments with moderate temporal uncertainty (i.e., $\sigma<0.03$). This trend is further supported by the additional metrics in Table I. Compared with ADG, WinkTPG generally achieves higher average velocity, lower total idle time, and fewer stop-and-go events under the Gaussian noise model, indicating smoother and more efficient motion. This supports our earlier claim that ADG is an overly conservative strategy. The solution quality of ADG remains nearly constant across different uncertainty levels because it replans at a very high frequency and is thus less sensitive to temporal uncertainty. As shown later, WinkTPG could achieve better solution quality with a smaller execution window.
@@ VII-B1 Unbounded Uncertainty Model
Effect of Window Size. Next, we evaluate the relationship between solution quality and the length of the execution and planning windows. All experiments are conducted in environments with $\sigma=0.03$. As shown in Fig. 9, configurations where the execution window is smaller than the planning window consistently show better solution quality. Also, smaller execution windows always achieve better performance, due to more frequent replanning with updated location information.
Figure 10: Empirical precedence violation rate and solution quality as a function of $P_{d}$ under the Gaussian noise model.
Effect of Safety Probability Threshold. We now examine how the theoretical safety probability threshold $P_{d}$ translates into empirical precedence violation rates under the Gaussian noise model. For each map and uncertainty level, we vary $P_{d}\in\{0.8,0.9,0.95,0.99\}$. Fig. 10 shows that the empirical violation rate remains below the theoretical upper bound $1-P_{d}$, confirming that the probabilistic safety margin behaves consistently with the Gaussian model assumptions, while the slight non-monotonicity on the random map is likely attributable to variance across trials. Meanwhile, larger $P_{d}$ values produce more conservative safety margins and therefore increased execution time, showing a clear trade-off between robustness and efficiency.
@@
TABLE II: Planning Runtime on the warehouse map with $\sigma=0.03$. MAPF refers to the time taken to generate the initial MAPF plan, while ADG and WinkTPG represent the runtime per planning round. For WinkTPG, we also report the actual average planning-window size $\bar{t_{p}}$, which denotes the average number of vertices included in the planning window during replanning.
Number of Agents MAPF ADG WinkTPG
100 $89.08\pm 32.76\,ms$ $0.24\pm 0.03\,ms$ $t_{p}$=10 $2.17\pm 0.20\,ms$ ($\bar{t_{p}}=11.30\pm 0.09$)
$t_{p}$=20 $3.31\pm 0.43\,ms$ ($\bar{t_{p}}=21.31\pm 0.10$)
$t_{p}$=40 $4.96\pm 0.78\,ms$ ($\bar{t_{p}}=41.34\pm 0.11$)
$t_{p}$$\infty$ $266.06\pm 52.33\,ms$ ($\bar{t_{p}}=N/A$)
1,000 $44513.62\pm 3941.23\,ms$ $2.02\pm 0.09\,ms$ $t_{p}$=10 $343.85\pm 14.32\,ms$ ($\bar{t_{p}}=12.62\pm 0.08$)
$t_{p}$=20 $389.23\pm 14.21\,ms$ ($\bar{t_{p}}=22.68\pm 0.08$)
$t_{p}$=40 $519.42\pm 20.13\,ms$ ($\bar{t_{p}}=42.82\pm 0.09$)
$t_{p}$$\infty$ $28211.74\pm 1673.22\,ms$ ($\bar{t_{p}}=N/A$)
TABLE III: Communication overhead on the warehouse map ($t_{e}=10$, $t_{p}=20$). Uplink denotes the total state updates sent from agents to the server, downlink denotes the total action commands sent from the server to agents, and replan freq. denotes the number of replanning cycles triggered per second.
Number of Agents Method Uplink (KB/replan) Downlink (KB/replan) Replan Freq. (Hz)
100 ADG $1.333\pm 0.111$ $2.134\pm 0.183$ $1.0$
WinkTPG $1.419\pm 0.099$ $19.953\pm 1.384$ $0.1$
1,000 ADG $13.474\pm 0.361$ $15.688\pm 0.461$ $1.0$
WinkTPG $13.258\pm 0.339$ $178.971\pm 4.901$ $0.1$
@@
Planning Runtime and Communication Overhead. We report in Table II the runtime of ADG, WinkTPG, and the MAPF planner. In both cases, ADG and WinkTPG demonstrate significantly faster runtimes compared to the MAPF planner used to generate the paths. Notably, while ADG appears faster, it requires significantly more frequent replanning. Additionally, with an appropriately chosen planning window size, WinkTPG generates plans in under a second, showing it is capable of near real-time performance. We further quantify the communication overhead in Table III. Although WinkTPG transmits larger messages per replanning cycle due to coordinated speed profiles, it replans less frequently than ADG. At 1,000 agents, the downlink per plan is approximately 179   KB, which remains well within the capacity of standard wireless networks.
VII-B2 Bounded Uncertainty Models
We next evaluate WinkTPG under the two bounded uncertainty models introduced in Sec. IV.
Unlike the Gaussian noise model, both bounded models admit deterministic safety guarantees, since the cumulative perturbations are bounded by construction.
Figure 11: Comparison of solution quality between ADG (solid lines) and WinkTPG (dashed lines) under the bounded cumulative delay model across different cumulative delay budget $\mathcal{K}$ (indicated by color).
Results under Bounded Cumulative Delay Model. For the bounded cumulative delay model, we vary the cumulative delay budget $\mathcal{K}\in\{1.0,2.0,3.0,5.0,10.0\}\,s$ and fix the unit delay $\eta=1.0\,s$. This parameter represents the maximum total execution delay (in seconds) an agent may accumulate along its path. As shown in  Figs. 11 and I, WinkTPG achieves 100% precedence satisfaction across all tested $\mathcal{K}$ values, confirming the deterministic guarantee derived in  Section IV. In terms of solution quality, WinkTPG consistently achieves lower suboptimality than ADG with smaller $\mathcal{K}$ values. The performance gap becomes more pronounced for moderate delay budgets (e.g., $\mathcal{K}=1.0$ and $\mathcal{K}=2.0$), where ADG’s conservative, event-driven execution causes unnecessary deceleration. As $\mathcal{K}$ increases, the solution cost increases gradually due to larger safety margins required to accommodate worst-case cumulative delays.
Figure 12: Comparison of solution quality between ADG (solid lines) and WinkTPG (dashed lines) under the bounded uniform noise model across different timing bound $\epsilon$ (indicated by color).
Results under Bounded Uniform Noise Model. For the bounded uniform noise model, we vary the timing bound $\epsilon\in\{0.01,0.02,0.1,0.2\}\,s$. Fig. 12 and Table I show the results. Again, WinkTPG achieves deterministic precedence satisfaction for all $\epsilon$ values tested. As expected, larger $\epsilon$ values lead to increased solution cost due to expanded safety intervals. Nevertheless, WinkTPG maintains better coordination than ADG, particularly for moderate uncertainty levels (e.g., $\epsilon=0.02$ and $\epsilon=0.1$), where proactive interval reasoning avoids excessive waiting.
Overall, the results in this subsection show that WinkTPG can effectively account for execution timing uncertainty and provide the intended robustness guarantees in practice. Under bounded uncertainty models, it achieves deterministic safety guarantees, while under the Gaussian noise model, the empirical violation rate remains below the theoretical bound. Moreover, WinkTPG performs well under moderate uncertainty levels and exhibits a clear efficiency-robustness tradeoff as uncertainty increases. As discussed in Section IV, the appropriate timing uncertainty model depends on the underlying controller and operating environment. For example, bounded models are suitable for settings where execution timing errors remain within a bounded range.
Figure 13: Simulation environments in Isaac-Sim.
VII-C Validation in Realistic Execution Settings
Finally, we validate WinkTPG under high-fidelity physics-based simulation and real-world experiments. In both settings, we use WinkTPG with the Gaussian noise model.
VII-C1 Physics-Based Simulation
We first validate WinkTPG using NVIDIA Isaac Sim [22]. This setting explicitly models robot dynamics, control latency, and contact interactions, allowing us to assess whether the speed profiles produced by kTPGu can be reliably executed by realistic robot platforms. We evaluate this setting on an Ubuntu 22.04 desktop machine with AMD 5800x processor and NVIDIA RTX 2080ti GPU.
Experiment Setup
We evaluate our method in two simulated environments: empty-small (shown in Fig. 13(a), size $8\times 8$) and random-small (shown in Fig. 13(b), size $16\times 16$), both converted from 2-D grid maps. In both environments, each grid cell has a physical size of $0.5\times 0.5\,\text{m}$. We use a differential-drive JetBot robot model with a diameter of $0.3\,\text{m}$. Each robot is subject to a linear speed limit of $0,1]\,\text{m/s}$, a linear acceleration limit of $0.5,0.5]\,\text{m/s}^{2}$, and an angular velocity limit of $0,\pi/4]\,\text{rad/s}$. A simple feedforward–feedback velocity controller is used to execute speed profiles produced by kTPGu. At each control step, the feedforward term computes the reference linear and angular velocities corresponding to the current time from the planned speed profile, while the feedback term computes the pose-tracking error between the reference pose implied by the planned trajectory and the robot’s actual pose. The control command is given by $u_{i}(t)=u_{i}^{\mathrm{ff}}(t)+K\,e_{i}(t)$, where $u_{i}^{\mathrm{ff}}(t)$ denotes the feedforward velocity command, $e_{i}(t)$ denotes the pose-tracking error, and K is a feedback gain matrix. The resulting control commands are applied via Isaac Sim’s articulation interface. To quantify tracking performance, we measure the timing error at the waypoints, $\Delta t_{i}(q)=\bigl|t_{i}^{\text{sim}}(q)-t_{i}^{r}(q)\bigr|$, where $t_{i}^{r}(q)$ is the planned arrival time determined by the speed profile and $t_{i}^{\text{sim}}(q)$ is the actual arrival time observed during execution. Empirically, the mean timing error is around 0.25 s, with maximum errors around 0.5 s. We use only simple parameter tuning for the controller; more advanced control strategies could further reduce tracking error but are beyond the scope of this work. We use 20 and 30 agents in the empty-small and random-small environments, respectively. For each environment, we evaluate 15 different MAPF plans generated by PBS under progressively increased uncertainty levels, with $\sigma\in\{0.0,0.001,0.005,0.01,0.05\}$. A run is considered successful if all agents reach their goals without any collision and no precedence violation happens. We report (1) the success rate over all trials, (2) and the sum of execution time over successful runs.
Figure 14: Sum of execution time and success rate as a function of uncertainty level $\sigma$ in physics-based simulation.
Experimental Results
As shown in Fig. 14, in both environments the success rate increases monotonically with the modeled uncertainty level. Higher uncertainty induces more conservative speed profiles and larger safety margins, thereby improving robustness to execution variability in physics-based simulation. This improved robustness comes at the cost of solution quality: as the uncertainty level increases, execution time grows noticeably, revealing a fundamental trade-off between robustness and efficiency. Consequently, the uncertainty level should be selected based on the target application and controller characteristics. Larger margins are appropriate for safety-critical scenarios or controllers with high variance, whereas smaller margins may suffice when higher efficiency is desired and execution can be tracked accurately.
Figure 15: Real-world robot experiments. The red line denotes the success rate, and the blue line denotes the sum of execution time.
VII-C2 Physical Robot Experiments
Finally, we demonstrate the execution of WinkTPG on real-world mobile robots.
Experiment Setup
We deploy WinkTPG on a team of 6 JetBot robots, consistent with the setup described in Section VII-C1, operating in a structured indoor environment as shown in Fig. 15. The workspace is configured as a $4\times 6$ grid, where each cell has a size of $0.3\,\text{m}\times 0.3\,\text{m}$. Both obstacles and workspace boundaries are physically defined. Robots communicate with a centralized coordination server over a local wireless network. A Vicon motion capture system provides real-time localization for all robots during execution. Similarly, we evaluate 5 different MAPF plans generated by PBS with progressively increasing uncertainty levels, with $\sigma\in\{0.0,0.01,0.02,0.05\}$.
Experimental Results
As shown in Fig. 15, WinkTPG executes reliably on physical robots under realistic sensing, actuation, and communication conditions. Increasing $\sigma$ leads to larger safety margins and more conservative speed profiles, resulting in longer execution times but improved reliability. The success rate stabilizes at 100% once sufficient margin is introduced, confirming that the proposed safety mechanism effectively compensates for execution uncertainty. These results demonstrate that the framework transfers from simulation to physical robots and exhibits the expected robustness-efficiency trade-off in practice.
VIII Conclusion